# Fast training and sampling of Restricted Boltzmann Machines

### v2

@@ Appendix B Sampling via Parallel Tempering using the learning trajectory
where $\Delta\mathcal{H}_{t}(\bm{x})=\mathcal{H}_{t}(\bm{x})-\mathcal{H}_{t-1}(\bm{x})$.
We continue alternating between the update step and the swap step until a total of $N_{\mathrm{mcmc}}$ steps is reached. The sampling procedure is illustrated in the following pseudo-code:
Input: Set of models $\{\mathcal{H}_{t}\}$, $t=0,\dots,t_{\mathrm{f}}$, Number of Gibbs steps $k$, Number of MCMC steps $N_{\mathrm{mcmc}}$
Output: Configurations $\bm{x}_{t}$ for $t=1,\ldots,t_{\mathrm{f}}$
Initialize: Sample $N$ chains from the RCM model $\bm{x}_{0}^{(0)}\sim\mathrm{RCMSampling}(\mathcal{H}_{0})$
for $t=1$ to $t_{\mathrm{f}}$ do
     $\bm{x}_{t}^{(0)}\sim\mathrm{GibbsSampling}(\mathcal{H}_{t},\bm{x}_{t-1}^{(0)},\tilde{k})$
@@
for $n=1$ to $\lfloor N_{\mathrm{mcmc}}/k\rfloor$ do
     for $t=1$ to $t_{\mathrm{f}}$ do
         $\bm{x}_{t}^{(n)}\sim\mathrm{GibbsSampling}(\mathcal{H}_{t},\bm{x}_{t}^{(n-1)},k)$
@@
     Resample $\bm{x}_{0}^{(n)}\sim\mathrm{RCMSampling}(\mathcal{H}_{0})$
     for $t=1$ to $t_{\mathrm{f}}$ do
         Compute acceptance probability
$p_{\mathrm{acc}}(\bm{x}_{t}^{(n)}\leftrightarrow\bm{x}_{t-1}^{(n)})=\min\left(1,\exp\left(\Delta\mathcal{H}_{t}(\bm{x}_{t}^{(n)})-\Delta\mathcal{H}_{t}(\bm{x}_{t-1}^{(n)})\right)\right)$
         Swap $\bm{x}_{t}^{(n)}$ and $\bm{x}_{t-1}^{(n)}$ with probability $p_{\mathrm{acc}}(\bm{x}_{t}^{(n)}\leftrightarrow\bm{x}_{t-1}^{(n)})$
@@
B.1 Controlling equilibration times with PTT
One of the advantages of PT-like approaches is that they allow easy control over thermalization times by studying the random walk in temperatures (or models for the PTT) visited by each run over time, as proposed in Ref. Banos et al. (2010). From now on, we will refer only to the number of models used in the PTT sampling, denoted as $N_{\mathrm{m}}$. In an ergodic sampling process, each run should spend roughly the same amount of time at each of the $N_{\mathrm{m}}$ different model indices, as all models have the same probability.
The mean temperature/model index is known to be $\left\langle n\right\rangle=(N_{\mathrm{m}}-1)/2$ (assuming the model indices run from 00 to $N_{\mathrm{m}}-1$). This simplifies the computation of the time-autocorrelation function of the model indices $n_{i}(t)$ visited by each sample $i$ at time $t$:
@@ B.1 Controlling equilibration times with PTT
as well as the integrated autocorrelation time, $\tau_{\mathrm{int}}$, using the self-consistent equation Sokal (1997); Amit & Martin-Mayor (2005):
@@
These two times provide different insights: $\tau_{\mathrm{exp}}$ gives insight into the time needed to thermalize the simulations (setting the length of the runs to at least $20\tau_{\mathrm{exp}}$ is generally a safe (and very conservative) choice to ensure thermalization Sokal (1997)), while $\tau_{\mathrm{int}}$ gives the time needed to generate independent samples. To keep track of the fluctuations, we compute a different $C^{m}(t)$ for each $m=1,\ldots,N_{\mathrm{m}}$ system index simulated in parallel in the PTT, where each $C^{m}(t)$ is averaged over 1000 independent realizations (with all runs initialized to the same model index). We then compute the average over all parallel runs $C(t)=\overline{C^{m}(t)}$ and use this curve to extract $\tau_{\mathrm{exp}}$ and $\tau_{\mathrm{int}}$. The errors in $C$ and the times are then extracted from the error of the mean of the values extracted with each of the $C^{m}(t)$ curves.
@@ Appendix D Computation of the likelihood using AIS
To compute the likelihood of the RBMs we used the Annealed Importance Sampling method introduced by Neal (2001) and recently studied in the case of the RBM Krause et al. (2020).
@@ D.1 Offline AIS
We follow the method described in Krause et al. (2020). The AIS estimate is based on creating a sequence of machines, at one end a machine from which we can compute the partition function and at the other end the machine from which we want to compute the partition function. Let’s consider that we have $i=0,\dots,N_{\rm rbm}$ machines described by
@@
where $p_{0}=p_{\rm ref}$ denote the RBM from which we can compute the partition function (typically a RBM where the weight matrix is zero) and $p_{N_{\rm rbm}}=p_{\rm target}$ the machine from which we want to estimate the partition function. For each machine, we denote its set of parameters $\bm{\Theta}_{i}=\{\bm{w}_{i},\bm{\theta}_{i},\bm{\eta}_{i}\}$. Following Krause et al. (2020), defining a transition rule $T_{i}(\bm{x}_{i-1}\to\bm{x}_{i})$ in order to pass from a configuration $\bm{x}_{i-1}$ of the system $i-1$ to $\bm{x}_{i}$ of the system $i$, that we have the following identity
@@ Appendix E Training of the RBM using the Jarzynski equation
In this section, we describe the procedure introduced in Carbone et al. (2024) for training energy-based models by leveraging the Jarzynki equation, and we adapt it to the specific case of the RBM. At variance with Carbone et al. (2024), we introduce an additional assumption on the transition probability of the parameters that will lead us to the the reweighting implemented in population annealing methods Weigel et al. (2021).
@@
Notice that, since Eq. equation 30 is an exact result, the importance weights should, in principle, eliminate the bias brought by the non-convergent chains used for approximating the log-likelihood gradient in the classical PCD scheme. However, after many updates of the importance weights, one finds that only a few chains carry almost all the importance mass. In other words, the vast majority of the chains we are simulating are statistically irrelevant, and we expect to get large fluctuations in the estimate of the gradient because of the small effective number of chains contributing to the statistical average. A good observable for monitoring this effect is the Effective Sample Size (ESS), defined as Carbone et al. (2024)
@@ Appendix F Tuning of the hyperparameters
In order to justify our choice of hyperparameters in the main text, we illustrate on Fig. 7 the value of the log-likelihood as a function of the number of hidden nodes and of the learning of the machine. In order to be fair, the comparison is done at a constant number of gradient updates of $10000$. In this regime, the machines that perform the best have a learning rate of $\gamma\sim 0.1,0.005$ and about $N_{h}=500,1000$ hidden nodes. We can expect that smaller learning rates reach the same value of the log-likelihood but for a much longer training. We clearly see the benefit of the RCM, where the log-likelihood achieves much higher values in general.
@@ Appendix G Continuous phase transition
where $g(E)$ the number of states with a given energy $E$, and thus $S(E)=\log g(E)$ is the entropy. Finally, the free energy is defined as $F(E)=E-S(E)$ and $f(E)=F(E)/N$ is the free energy per variable. All $Z$, $E$ and $F$ depend on the model parameters. $N$ refers to the number of variables.
@@
In the case of the RBM, the first learning transition can be mapped to a very simple model, the so-called Curie-Weiss model, as recently described in Ref. Bachtis et al. (2024). The Curie-Weiss model is fully solvable, and for this reason the meaning of a continuous transition is perhaps easier to understand there. Let’s give it a try. The Curie-Weiss model consists in a set of $N$ discrete variables $s_{i}=\pm 1$. We define the (exponential) distribution over these variables as
@@
we expect that for $\beta$ small each spin behaves as an isolated Rademacher random variables (each spin being ±$\pm 1$ with probability one-half), while for large value of $\beta$, the distribution of the system is dominated by configurations where the variables have (almost) all the same signs. The key mathematical aspect is to study the system in the limit where $N\to\infty$. A way to study this distribution is to investigate the moment generating function. It is possible to show (rigorously in this precise model), that the probability of $m=N^{-1}\sum_{i}s_{i}$, also called the magnetization of the system, is given by $p(m)\propto\exp(-N\Omega(m))$ (in the large $N$ limit) where $\Omega(m)$ is a large deviation function. In the CW model, it is given by
@@
The structure of $\Omega(m)$ is such that for $\beta<\beta_{c}=1$, it is a convex function with only one minimum in $m=0$, while for $\beta>\beta_{c}$ it has two symmetric minima $m=\pm m(\beta)\neq 0$ with $\Omega(m(\beta))=\Omega(-m(\beta))$ (the two minimum are equally probable). The values (in function of $\beta$) of the $m$ are continuous: $m=0$ for $\beta\in[0,1]$, and then positive until it saturates to $1$ as $\beta\to\infty$. In practice, it means that the distribution $p(m)$ pass from an unimodal distribution to a bimodal one as $\beta$ increases. The point where it happens is at $\beta_{c}=1$ and is called a second order (continuous) phase transition because de value of $m$ changes continuously from zero to non-zero values. A particular interesting properties is that at $\beta_{c}$, the function $\Omega(m)$ develops a non-analyticity in its second-order derivative, that can be linked with long-range fluctuations. For a recent review Kochmański et al. (2013), and for more rigorous results Sinai (2014).
@@
If $H>H_{\rm sp}(\beta)$, then the large deviation function has only one minimum for positive $m$. Now if $H<H_{\rm sp}(\beta)$, then $\Omega$ will present two minima. A first one for positive $m$ and a second one for negative $m$. The reason is that the pairwise interactions “manage” to create a metastable state for negative magnetization even in the presence of a positive field or bias. It is clear that this state will be unstable w.r.t. the one with positive magnetization which traduces in a minimum of higher value in the large deviation function. By changing the value of $H$ from large positive values toward large negative ones, we then pass from a system with only one state, then two states with one being subdominant, and finally back to one state but where the magnetization has changed its sign. In the infinite size limit, the system will always dominated by the state where $\Omega$ is at its lowest value. However, if one prepare a system close to a metastable state, the system will remain trap there until it disappears.
@@
First, it has been shown in previous works Decelle et al. (2018); Bachtis et al. (2024) that the RBM’s training undergoes a set of continuous transitions, at least at the beginning of the learning. Hence, changing the parameters of the RBM following the learning trajectories, the transitions are continuous. Then in Decelle & Furtlehner (2021a) it is shown that the relaxed version of the RBM, namely the RCM undergoes a first-order transition when changing the temperature. Schematically to understand that, it is enough to see how the parameters $\beta$ acts on the system. In the case of the RCM, the energy term behaves in the standard way, i.e. linearly with the inverse temperature $\beta$ and the free energy takes the ordinary form
@@
Appendix H Overfitting and privacy loss as quality indicators

### v3

@@ Appendix B Sampling via Parallel Tempering using the learning trajectory
where $\Delta\mathcal{H}_{t}(\bm{x})=\mathcal{H}_{t}(\bm{x})-\mathcal{H}_{t-1}(\bm{x})$. We continue alternating between the update step and the swap step until a total of $N_{\mathrm{mcmc}}$ steps is reached.
B.1 Pseudo-code of PTT vs PT
We provide here the pseudo code for both Parallel Tempering (Alg. 1) and Parallel Trajectory Tempering (Alg. 2) sampling schemes. The difference between both algorithms lies in the way the parallel models are selected. In the case of Parallel Tempering, an annealing on the temperature is made, whereas for Parallel Trajectory Tempering several models are saved during training.
Algorithm 1 Parallel Tempering
$\{\beta_{j}\}_{j=\{1,\ldots,\texttt{N}_{\beta}\}},\bm{\Theta}=\{\bm{w},\bm{\theta},\bm{\eta}\},\texttt{N}_{\texttt{iter}},\texttt{N}_{\texttt{increment}},\texttt{N}_{\texttt{sample}}$
for $j=1,\ldots,N_{\beta}$ do
  $\texttt{chains}[j]=\texttt{RandomBinaryChains}(\texttt{N}_{\texttt{sample}})$
@@ B.1 Pseudo-code of PTT vs PT
for $i=1,\ldots,\texttt{N}_{\texttt{iter}}$ do
  for $j=1,\ldots,\texttt{N}_{\beta}$ do
   $\texttt{chains}[j]\leftarrow\texttt{AlternateGibbsSampling}(\texttt{N}_{\texttt{increment}},\beta_{j}\bm{\Theta},\texttt{chains}[j])$
@@
  for $j=1,\ldots,\texttt{N}_{\beta}-1$ do
   $\texttt{chains}[j],\texttt{chains}[j+1]\leftarrow\texttt{SwapConfigurations}(\texttt{chains}[j],\beta_{j}\bm{\Theta},\texttt{chains}[j+1],\beta_{j+1}\bm{\Theta})$
@@
Algorithm 2 Parallel Trajectory Tempering
$\{\bm{\Theta}_{j}=\{\bm{w}_{j},\bm{\theta}_{j},\bm{\eta}_{j}\}\}_{j=\{1,\ldots,\texttt{N}_{\texttt{model}}\}},\texttt{N}_{\texttt{iter}},\texttt{N}_{\texttt{increment}},\texttt{N}_{\texttt{sample}}$
for $j=1,\ldots,\texttt{N}_{\texttt{model}}$ do
  $\texttt{chains}[j]=\texttt{RandomBinaryChains}(\texttt{N}_{\texttt{sample}})$
end for
for $i=1,\ldots,\texttt{N}_{\texttt{iter}}$ do
  for $j=1,\ldots,\texttt{N}_{\texttt{model}}$ do
   $\texttt{chains}[j]\leftarrow\texttt{AlternateGibbsSampling}(\texttt{N}_{\texttt{increment}},\bm{\Theta}_{j},\texttt{chains}[j])$
  end for
  for $j=1,\ldots,\texttt{N}_{\texttt{model}}-1$ do
   $\texttt{chains}[j],\texttt{chains}[j+1]\leftarrow\texttt{SwapConfigurations}(\texttt{chains}[j],\bm{\Theta}_{j},\texttt{chains}[j+1],\bm{\Theta}_{j+1})$
  end for
end for
B.2 Controlling equilibration times with PTT
One of the advantages of PT-like approaches is that they allow easy control over thermalization times by studying the random walk in temperatures (or models for the PTT) visited by each run over time, as proposed in Ref. (banos2010nature). From now on, we will refer only to the number of models used in the PTT sampling, denoted as $N_{\mathrm{m}}$. In an ergodic sampling process, each run should spend roughly the same amount of time at each of the $N_{\mathrm{m}}$ different model indices, as all models have the same probability.
The mean temperature/model index is known to be $\left\langle n\right\rangle=(N_{\mathrm{m}}-1)/2$ (assuming the model indices run from 0 to $N_{\mathrm{m}}-1$). This simplifies the computation of the time-autocorrelation function of the model indices $n_{i}(t)$ visited by each sample $i$ at time $t$:
@@ B.2 Controlling equilibration times with PTT
as well as the integrated autocorrelation time, $\tau_{\mathrm{int}}$, using the self-consistent equation (sokal1997monte; amit2005field):
@@
These two times provide different insights: $\tau_{\mathrm{exp}}$ gives insight into the time needed to thermalize the simulations (setting the length of the runs to at least $20\tau_{\mathrm{exp}}$ is generally a safe (and very conservative) choice to ensure thermalization (sokal1997monte)), while $\tau_{\mathrm{int}}$ gives the time needed to generate independent samples. To keep track of the fluctuations, we compute a different $C^{m}(t)$ for each $m=1,\ldots,N_{\mathrm{m}}$ system index simulated in parallel in the PTT, where each $C^{m}(t)$ is averaged over 1000 independent realizations (with all runs initialized to the same model index). We then compute the average over all parallel runs $C(t)=\overline{C^{m}(t)}$ and use this curve to extract $\tau_{\mathrm{exp}}$ and $\tau_{\mathrm{int}}$. The errors in $C$ and the times are then extracted from the error of the mean of the values extracted with each of the $C^{m}(t)$ curves.
@@ Appendix D Computation of the likelihood using AIS
To compute the likelihood of the RBMs we used the Annealed Importance Sampling method introduced by (neal2001annealed) and recently studied in the case of the RBM (krause_algorithms_2020).
@@ D.1 Offline AIS
We follow the method described in (krause_algorithms_2020). The AIS estimate is based on creating a sequence of machines, at one end a machine from which we can compute the partition function and at the other end the machine from which we want to compute the partition function. Let’s consider that we have $i=0,\dots,N_{\rm rbm}$ machines described by
@@
where $p_{0}=p_{\rm ref}$ denote the RBM from which we can compute the partition function (typically a RBM where the weight matrix is zero) and $p_{N_{\rm rbm}}=p_{\rm target}$ the machine from which we want to estimate the partition function. For each machine, we denote its set of parameters $\bm{\Theta}_{i}=\{\bm{w}_{i},\bm{\theta}_{i},\bm{\eta}_{i}\}$. Following (krause_algorithms_2020), defining a transition rule $T_{i}(\bm{x}_{i-1}\to\bm{x}_{i})$ in order to pass from a configuration $\bm{x}_{i-1}$ of the system i1i-1 to $\bm{x}_{i}$ of the system $i$, that we have the following identity
@@ Appendix E Training of the RBM using the Jarzynski equation
In this section, we describe the procedure introduced in (carbone2024efficient) for training energy-based models by leveraging the Jarzynki equation, and we adapt it to the specific case of the RBM. At variance with (carbone2024efficient), we introduce an additional assumption on the transition probability of the parameters that will lead us to the the reweighting implemented in population annealing methods (PhysRevE.103.053301).
@@
Notice that, since Eq. equation 30 is an exact result, the importance weights should, in principle, eliminate the bias brought by the non-convergent chains used for approximating the log-likelihood gradient in the classical PCD scheme. However, after many updates of the importance weights, one finds that only a few chains carry almost all the importance mass. In other words, the vast majority of the chains we are simulating are statistically irrelevant, and we expect to get large fluctuations in the estimate of the gradient because of the small effective number of chains contributing to the statistical average. A good observable for monitoring this effect is the Effective Sample Size (ESS), defined as (carbone2024efficient)
@@ Appendix F Tuning of the hyperparameters
In order to justify our choice of hyperparameters in the main text, we illustrate on Fig. 7 the value of the log-likelihood as a function of the number of hidden nodes and of the learning of the machine. In order to be fair, the comparison is done at a constant number of gradient updates of $10000$. In this regime, the machines that perform the best have a learning rate of $\gamma\sim 0.1,0.005$ and about ${N_{\mathrm{h}}}=500,1000$ hidden nodes. We can expect that smaller learning rates reach the same value of the log-likelihood but for a much longer training. We clearly see the benefit of the RCM, where the log-likelihood achieves much higher values in general.
@@ Appendix G Continuous phase transition
where $g(E)$ the number of states with a given energy $E$, and thus $S(E)=\log g(E)$ is the entropy. Finally, the free energy is defined as F(E)=ES(E)F(E)=E-S(E) and $f(E)=F(E)/N$ is the free energy per variable. All $Z$, $E$ and $F$ depend on the model parameters. $N$ refers to the number of variables.
@@
In the case of the RBM, the first learning transition can be mapped to a very simple model, the so-called Curie-Weiss model, as recently described in Ref. (bachtis2024cascade). The Curie-Weiss model is fully solvable, and for this reason the meaning of a continuous transition is perhaps easier to understand there. Let’s give it a try. The Curie-Weiss model consists in a set of $N$ discrete variables $s_{i}=\pm 1$. We define the (exponential) distribution over these variables as
@@
we expect that for $\beta$ small each spin behaves as an isolated Rademacher random variables (each spin being $\pm 1$ with probability one-half), while for large value of $\beta$, the distribution of the system is dominated by configurations where the variables have (almost) all the same signs. The key mathematical aspect is to study the system in the limit where $N\to\infty$. A way to study this distribution is to investigate the moment generating function. It is possible to show (rigorously in this precise model), that the probability of $m=N^{-1}\sum_{i}s_{i}$, also called the magnetization of the system, is given by $p(m)\propto\exp(-N\Omega(m))$ (in the large $N$ limit) where $\Omega(m)$ is a large deviation function. In the CW model, it is given by
@@
The structure of $\Omega(m)$ is such that for $\beta<\beta_{c}=1$, it is a convex function with only one minimum in $m=0$, while for $\beta>\beta_{c}$ it has two symmetric minima $m=\pm m(\beta)\neq 0$ with $\Omega(m(\beta))=\Omega(-m(\beta))$ (the two minimum are equally probable). The values (in function of $\beta$) of the $m$ are continuous: $m=0$ for $\beta\in[0,1]$, and then positive until it saturates to $1$ as $\beta\to\infty$. In practice, it means that the distribution $p(m)$ pass from an unimodal distribution to a bimodal one as $\beta$ increases. The point where it happens is at $\beta_{c}=1$ and is called a second order (continuous) phase transition because de value of $m$ changes continuously from zero to non-zero values. A particular interesting properties is that at $\beta_{c}$, the function $\Omega(m)$ develops a non-analyticity in its second-order derivative, that can be linked with long-range fluctuations. For a recent review (kochmanski2013curie), and for more rigorous results (sinai2014theory).
@@
If $H>H_{\rm sp}(\beta)$, then the large deviation function has only one minimum for positive $m$. Now if $H<H_{\rm sp}(\beta)$, then $\Omega$ will present two minima. A first one for positive $m$ and a second one for negative $m$. The reason is that the pairwise interactions “manage” to create a metastable state for negative magnetization even in the presence of a positive field or bias. It is clear that this state will be unstable w.r.t. the one with positive magnetization which traduces in a minimum of higher value in the large deviation function. By changing the value of $H$ from large positive values toward large negative ones, we then pass from a system with only one state, then two states with one being subdominant, and finally back to one state but where the magnetization has changed its sign. In the infinite size limit, the system will always dominated by the state where $\Omega$ is at its lowest value. However, if one prepares a system close to a metastable state, the system will remain trapped there until it disappears.
@@
First, it has been shown in previous works (decelle2018thermodynamics; bachtis2024cascade) that the RBM’s training undergoes a set of continuous transitions, at least at the beginning of the learning. Hence, changing the parameters of the RBM following the learning trajectories, the transitions are continuous. Then in (decelle2021exact) it is shown that the relaxed version of the RBM, namely the RCM undergoes a first-order transition when changing the temperature. Schematically to understand that, it is enough to see how the parameters $\beta$ acts on the system. In the case of the RCM, the energy term behaves in the standard way, i.e. linearly with the inverse temperature $\beta$ and the free energy takes the ordinary form
@@
Appendix H First order transitions in RBMs
In this section, we examine the appearance of first-order transitions when adjusting the temperature of a well-trained RBM on clustered datasets—a phenomenon that significantly hinders the performance of thermal sampling algorithms. We first illustrate this issue using an analytically tractable toy model, followed by an analysis of the first-order temperature transition observed in the RBM trained on the HGD dataset, as discussed in the main text.
H.1 Theoretical analysis in a simple model
We propose a simple model of dataset on which we will show that the learned RBM is having a first order transition in temperature. We first consider an artificial dataset using a Curie-Weiss model. We defined the following Hamiltonian on the variables $s_{i}=\pm 1$
$\mathcal{H}=-\frac{1}{2{N_{\mathrm{v}}}}\left(\sum_{i}s_{i}\right)^{2},\quad p(\bm{s})=\frac{1}{Z}\exp(-\beta_{T}\mathcal{H}).$
When the inverse temperature $\beta_{T}$ is below one, the distribution in the $N_{\mathrm{v}}$$\to\infty$ limit is unimodal, with the Boltzmann average of the variables $s$ being $\langle s\rangle=0$. For $\beta_{T}>1$, the system becomes bimodal, with $\langle s_{i}\rangle=m$, where $m$ satisfies the self-consistent equation $m=\tanh(\beta_{T}m)$. We propose to model this dataset using a Bernoulli-Gauss RBM with one hidden node. The visible nodes, $\sigma=\{0,1\}$, follow a Bernoulli distribution, while the hidden node, $\tau$, follows a Gaussian distribution with zero mean and variance $1/{N_{\mathrm{v}}}$, combined with a local visible bias. When the inverse temperature $\beta_{T}$ is below one, the distribution in the $N_{\mathrm{v}}$$\to\infty$ limit is unimodal and the average w.r.t. the Boltzmann distribution of the values of the variables $s$’s is $\langle s\rangle=0$. When $\beta_{T}$ is above one, the system becomes bimodal with $\langle|s_{i}|\rangle=m$ when $m$ is defined by the self-consistent equation $m=\tanh(\beta_{T}m)$. We now propose to learn this dataset using a Bernoulli-Gauss RBM with one hidden node, where the visible nodes are $\sigma=\{0,1\}$, and the hidden one $\tau$ is Gaussian distribution of zero mean and variance $1/{N_{\mathrm{v}}}$, and a local visible bias. The Hamiltonian is then
$\mathcal{H}=-\sum_{i}\sigma_{i}W_{i}\tau-\sum_{i}\eta_{i}\sigma_{i},$
where $W_{i}$ is the weight matrix and $\eta_{i}$ the biases. The leaning then consist in converting the binary variable of the Curie-Weiss model into $\{0,1\}$ and to learn the coresponding dataset. By a rapid inspection (in the infinite size limit), we can easily show that the optimal learned parameters for the RBM are
$w_{i}=2\sqrt{\beta_{T}}\text{ and }\eta_{i}=-2\beta_{T}.$
Now, thanks to the simplicity of the model, we can actually compute the free energy of the model where the Hamiltonian has been rescaled by an annealing temperature $\beta$, we obtain
$-f(m_{T})$ $=\frac{m_{\tau}^{2}}{2}-N^{-1}\sum_{i}\log\left[1+\exp(\beta w_{i}m_{\tau}+\beta\eta_{i})\right]$
$\text{ where }m_{\tau}=\frac{1}{N}\sum_{i}\beta w_{i}{\rm sigm}\left[1+\exp(\beta w_{i}m_{\tau}+\beta\eta_{i}\right]$
In Fig. 9, we plot $\beta f(m_{T})$ for ten different values of $\beta$ within the range $[0.8,1.05]$. It is evident that $\beta=1$ represents the coexistence point, indicating a first-order transition in temperature. Note that the probability $p(m)\approx e^{-N\beta f(m_{T})}$ for large $N$, implying that for $\beta>1$, the typical configuration corresponds to $m>1$, whereas for $\beta<1$, it corresponds to $m<1$. This clearly shows that even on simple case we observe such transition when doing annealing.
Figure 9: Free energy $\beta f(m_{\tau})$ for $\beta\in[0.8,1.05]$ for the optimal RBM’s parameters learned on a CW model with $\beta_{T}=1.4$. At $\beta=1.0$ the two minima are perfectly adjusted. As we change the temperature of the RBM, they get destabilized showing the presence of a first order transition.
H.2 First order phase transition in temperature on a trained RBM on a real dataset
We now consider the RBM trained on the HGD dataset using the pretraining and 50 000 updates of PCD training. As we vary the temperature, we recover the probability $p(\bm{m})$ learned by the model using the TMC method (bereux2023learning), projected along the second principal component in Fig. 11 and on both first and second principal components in Fig. 10. While $\beta<1.0$, the probability distribution is dominated by the central mode, equilibrating between all modes $\beta=1.0$. As $\beta$ goes above 1.0, the distribution becomes dominated by the two external clusters.
The effect is clearly visible when sampling the model using the PT algorithm (Fig. 12). The sampling was performed with 200 temperatures with a minimum acceptance rate of 0.78 between two temperatures. We see a dataset cluster being ignored with only a few samples in it. However, when resampling sampling the model for $10^{6}$ AGS steps starting from the samples obtained through PT, we see a clear change in the distribution, equilibrating with the missing cluster. In contrast, performing the same experiment with PTT instead of PT, we observe no shift in the empirical distribution of the samples after $10^{6}$ AGS, as the upper cluster is already sampled.
Figure 10: Projection of the negative log probability (top) and probability (bottom) learned by the model, constrained to visible magnetizations $m_{0}$ and $m_{1}$, for various values of $\beta$. The red dots represent the training data projected onto the first two principal components.
Figure 11: Negative log probability (left) and probability (right) learned by the model constrained to have a given $m_{1}$ for several values of $\beta$. The equilibrium distribution of the model is displayed as the black dotted line ($\beta=1$).
Figure 12: Comparison of the sampling results of PT (left) and PTT (right). The top row corresponds to sample generated with each methods compared with the dataset distribution. The bottom row compares samples obtained by performing $10^{6}$ AGS starting from the PT and PTT samples with their starting position.
Appendix I Overfitting and privacy loss as quality indicators